# EPOCHS IV: SED Modelling Assumptions and their impact on the Stellar Mass Function at using PEARLS and public JWST observations

### v1

@@ 4.2 Prospector
In addition to our Bagpipes SED-fitting we also fit our sample using the Prospector package (Johnson et al., 2021) in order to compare the results of these two commonly used SED fitting tools. Prospector allows greater flexibility and control of model parameters than Bagpipes at the expense of computational time. Prospector uses Bayesian inference to determine galaxy stellar population properties and star formation histories. Prospector is built on Flexible Stellar Population Synthesis (FSPS, Conroy & Gunn, 2010)), using python-fsps (Johnson et al., 2023) and the Modules for Experiments in Stellar Astrophysics Isochrones and Stellar Tracks (MIST) stellar isochrones (Choi et al., 2016; Dotter, 2016). Prospector allows for very flexible star formation histories, including the use of non-parametric SFHs, which have become increasingly popular in the JWST era. Non-parametric SFHs are modelled by splitting star-formation history into a number of temporal bins, with the SFH in each bin linked by a statistical distribution. We generally follow the prescription of Tacchella et al. (2022) for our Prospector model. We test both a traditional parametric star formation history as well as the “continuity bursty” non-parametric SFH used in Bagpipes. We allow the stellar mass to vary between 6 12 with a uniform logarithmic prior. For our parametric star formation history we use the ‘delayed-’ model, where .
@@ 4.3 Stellar Mass Comparisons
Figure 1 shows a comparison of the derived galaxy properties of our fiducial Bagpipes run for all sources in our sample to our other Bagpipes results. The details of the models compared are shown in Table 2. As discussed in § 4.1, we compare models with different priors, SPS models and SFH parametrizations. The top plot shows the systematic mass offset found when varying the dust law and priors, while the second row shows the same offset for parametric and non-parametric SFHs, as well as a BPASS SPS model (Stanway & Eldridge, 2018). For each Bagpipes model we show the Locally Weighted Scatterplot Smoothing fit (LOWESS, Cleveland, 1979). The LOWESS estimator is a non-parametric fitter for noisy data, and shows the overall trend between the mass discrepancy and fiducial stellar mass.
We allow the redshift to vary within the EAZY-py posterior PDF for each Bagpipes fit, resulting in only small changes in redshift between individual Bagpipes results. 99.6% of photo- estimates fall within 15% of our fiducial Bagpipes redshift.
In the following sections we compare the consistency of the derived stellar masses between our fiducial and alternative Bagpipes models, as shown in Figure 1. We also include some limited discussion of other inferred galaxy properties, such as star formation rate, or dust content where they vary significantly from our fiducial model.
In Figure 3 and Figure 4 we show examples of the photometry, best-fitting Bagpipes & Prospector SEDs, and posterior redshift and stellar mass estimates for a selection of high-mass and or discrepant galaxies in our sample.
4.3.1 Impact of Priors
The first comparison we make is a substitution of the default logarithmic prior on the V-band dust extinction, A, to a uniform prior. A uniform prior favors higher dust extinctions, and the largest difference to the fiducial model is seen in galaxies with A mag, where the dust content is poorly constrained. We see good agreement in photo-, along with a few extreme outliers, which is unsurprising given the informative redshift prior from EAZY-py we use. The stellar mass offset is shown in light blue in the top plot of Figure 1 and shows good agreement within the posterior uncertainties. Galaxies with higher levels of dust are found to be significantly younger and more star-forming with a uniform dust prior, suggesting the inferred star formation histories are dependent on the dust prior. Comparison of the best-fitting values shows that both models are equally well-fitted to the photometry.
We also test the impact of our metallicity prior, which is logarithmic in our fiducial Bagpipes model. This favours low metallicity, which we expect in the early Universe. Here we exchange this prior for a uniform distribution which favours higher metallicity, shown in orange in the top plot of Figure 1. We see little overall impact on the stellar mass from the metallicity prior, with individual galaxies scattering up to 0.5 dex and the majority consistent with results from our fiducial Bagpipes run.
4.3.2 The Assumed Dust Law
Our fiducial Bagpipes model assumes a simple one-component Calzetti et al. (2000) dust law. The slope of the dust law is known to vary in some galaxies, and numerous alternative models (e.g. Charlot & Fall, 2000; Salim et al., 2018) including additional parameters have been suggested. Charlot & Fall (2000) fit a two-component dust law, with different amounts of extinction for young ( Myr) and old stellar populations, to account for dust in stellar birth clouds. The resultant deviations in stellar mass are shown in magenta in the top plot of Figure 1, with a significant deviation at the highest stellar masses, which can be seen in the LOWESS trend. For galaxies with a fiducial stellar mass of M, the majority are found to have 1 dex larger stellar masses with the Charlot & Fall (2000) dust law. However for these galaxies the goodness of fit is considerably poorer, with with the Charlot & Fall (2000) model.
Despite the same prior, the dust attenuation (A) posteriors are quite different between the two models, due to the degeneracy between A and slope. Galaxies with moderate dust extinction in the fiducial results ( mag) are typically found to have very little dust extinction on the old stellar population, but the most dust-obscured galaxies (A) have significantly greater extinction (1.5 more on average). The majority of galaxies favor a steeper slope () for the attenuation power-law than given by Calzetti et al. (2000), in this mode with a distribution centered on n, which is 2 from the prior value of 0.70.
Salim et al. (2018) allow a deviation in slope compared to the Calzetti et al. (2000) and an additional UV bump at . The UV bump in the Salim et al. (2018) model is driven primarily by Polycyclic Aromatic Hydrocarbon (PAH) emission, which is expected when emission is strongly reprocessed. We replace the Calzetti et al. (2000) dust law with the Salim et al. (2018) model, and fit for these additional components, as shown in Table 2. We find that the inferred stellar mass can increase by 1 dex in some cases, particularly at the highest fiducial stellar mass, but the average offset (see the LOWESS fit in Figure 1 is smaller than for the Charlot & Fall (2000) case. In a number of cases the goodness of fit for the Salim et al. (2018) model is considerably better than our fiducial model, with in a number of galaxies.
4.3.3 Comparison to ‘delayed’ SFH
To test the consistency of galaxy stellar mass estimates with different parametric star formation histories, we replace our ‘lognormal’ SFH with another commonly used SFH; a delayed- SFH. We find systematically slightly lower stellar masses above 10 M, but reasonable agreement at the lowest stellar masses. This is likely due to the log-uniform prior on age used in the delayed- model, which results in a younger stellar population. Comparison of the goodness of fit via the parameter suggests the models typically have slightly poorer fits than the fiducial model, but in some cases the stellar masses are reduced by 0.5 dex with very little impact on the goodness of fit.
4.3.4 Comparison to “continuity bursty” SFH
Here we replace our fiducial Bagpipes ‘lognormal’ SFH with a “continuity bursty” SFH model described in detail in § 4.1. We reproduce the result of Tacchella et al. (2022) that this SFH increases the galaxy stellar mass, finding an average increase of 0.2 dex. The stellar mass discrepancy between the two models is shown in purple in the lower figure of 1. In individual cases the increase in stellar mass can reach 1 dex, with only small changes in photo-. The largest offsets in stellar mass are typically seen for galaxies with a fiducial stellar mass of M, with the highest fiducial stellar mass galaxies ( M) seeing considerably smaller increases.
4.3.5 Comparison to BPASS SPS Models
By default Bagpipes uses the 2016 version of Bruzual & Charlot (2003)’s (BC03) stellar population synthesis (SPS) models. However Bagpipes can also employ Binary Population and Spectral Synthesis (BPASS, Stanway & Eldridge, 2018) models. Specifically we use models generated with v2.2.1 of BPASS, with the default IMF (slope of 1.35, 300 M). The IMF parametrization of this SPS model differs slightly from the Kroupa (2001) IMF used in the BC03 models, causing an intrinsic offset in stellar mass. However as shown in orange in the lower plot of Figure 1, the comparison of the mass estimates is more complex. At the lowest and highest stellar masses ( and ), we see systematically lower stellar masses than the fiducial Bagpipes model, which is what we would expect based the IMF difference alone. However for the majority of the sample, which falls between these two mass extremes, we see a significantly larger offset in stellar masses, which increases with redshift, as shown by the LOWESS trend.
4.3.6 Other Comparisons
2. “continuity” non-parametric SFH model’: This variation tests the other non-parametric SFH model introduced in Leja et al. (2019) and Tacchella et al. (2022), which more tightly constrains the SFR in neighbouring time bins to force more smoothly varying star formation histories than the “continuity bursty” model. We find overall similar behaviour to the “continuity bursty” model, with systematically higher masses and higher mass-weighted age. The models typically have worse statistics than the “continuity bursty” model fits potentially indicating that more stochastic star formation histories are preferred for the majority of galaxies, as suggested in the literature (Faucher-Giguère, 2018; Looser et al., 2023; Asada et al., 2023; Endsley et al., 2023a).
4.4 Comparison to Prospector
For the comparison of the parametric SFH models (delayed- for Prospector, lognormal for Bagpipes), we see systematically larger photo- and stellar mass estimates with Prospector. Whilst in the Prospector model we allow the IGM attenuation to vary, which could impact photo- estimates, we do not see the same offset with the non-parametric SFH model, where the IGM attenuation is also allowed to vary. Both SED-fitting tools are given the same redshift prior from EAZY-py. We can see for both example galaxy SEDs in Figure 4 that Prospector instead prefers higher- solutions than Bagpipes because it is inferring the presence of Lyman- emission. Whilst Lyman- emitters have been found at (e.g. Bunker et al., 2023b), we do not expect to observe Lyman- from the majority of galaxies at these redshifts due to the attenuation from neutral hydrogen during the Epoch of Reionization, and a systematic photo- offset from Prospector is likely for galaxies without Lyman- emission. However given that we see a photo offset primarily for the ’parametric SFH’ only, Lyman- emission is unlikely to be the only cause of the offset as the Prospector SEDs in Figure 4 show Lyman- emission for both SFH models. This offset will generally cause a slight increase in stellar mass estimates, as a more distant galaxy must be intrinsically brighter. We see a median increase of 0.43 dex in stellar mass.
@@ 4.5 Top-heavy IMF
In Figure 5 we show the results of modifying the IMF on the stellar mass estimates. We compare the masses derived with a standard Kroupa (2001) IMF to the two modified IMF models, finding significant decreases in stellar mass with very little change in the quality of the fits. Noting that due to the computational intensity of Prospector fitting we only fit a subset of our full sample, and calculate the median decrease in mass for both SFH models and HOT IMFs. For the sample where we use the HOT 60K IMF we include our full parent sample, however for the model we fit only galaxies with either or fiducial Bagpipes stellar mass . In terms of numbers, there are 221 galaxies in the HOT 45K IMF group, and 21 galaxies in the HOT IMF 60K group. Given that this IMF model is predicted to be applicable in the region of 8 12 (Steinhardt et al., 2023), and the lack of mass-dependence in our results, we do not expect the excluded lower mass galaxies at to significantly impact our findings.
For the non-parametric “delayed” SFH model, we find a median decrease in stellar mass of 0.33 dex for the (HOT 45K IMF) galaxy sample. For the sample, we find a median decrease of 0.46 dex. For the “continuity bursty” SFH model, we find a median stellar mass decrease of 0.26 dex for the 12 sample (HOT 45K IMF) and a decrease of 0.36 dex for the 12 sample (HOT 60K IMF). As expected, we see a larger decrease in stellar mass for the HOT 60K model, which is more ‘top-heavy’ than the HOT 45K model.
@@ 5 Galaxy Stellar Mass Functions
We make a further cut to our galaxy sample here, removing the fields of El Gordo, SMACS-0723 and Clio. The depths and available filters of these datasets mean that they do not contribute significant volume to our GSMF estimates, but increase the redshift uncertainties (Adams et al., 2023). This reduces the number of galaxies used in the GSMF to 1092, and the total area used to 175 .
@@
where are the co-moving distances at the maximum/minimum detectable redshifts (capped at bin edges). are the survey areas, which are given in Table 1. For the purposes of constructing a GSMF we exclude the Clio, El Gordo and SMACS-0723 fields due to their shallow depth or lack of key broadband photometry (e.g. F115W).
@@ 5.1 Detection Completeness
The magnitude range was swept in steps of 0.2, with 1000 sources inserted into the image at a time in order to prevent introducing artificial overdensities. We assume a size - luminosity relation of (Grazian et al., 2012), with reference size set to pc at , the maximum available size was on the order of 1000 pc. The inclination angle of these sources was assumed to be uniformly random. We sweep a redshift range of 6.5 to 13.5 in steps of 0.5 to account for the changes in angular diameter distance and obtain a 2D dependence of completeness on redshift and absolute magnitude.
@@ 5.2 Selection Completeness and Contamination
We generate mock observations from these catalogues by estimating 1 errors in the measurements based on the average depths of our JWST reductions in each field (see Table 1). We then perturb the catalogue photometry in flux space within a Gaussian centered on the catalogue measurement and width equal to the 1 error in that filter. We then run the galaxies through our full selection procedure to measure how well our pipeline recovers the true redshifts of the sources. This allows us to compare the robustness of our selection criteria given the differing depths and filters in each set of observations.
@@ 5.2.3 Active Galactic Nuclei
Given the apparent prevalence of high- AGN detected by multi-wavelength and spectroscopic studies it is likely there is a population of galaxies containing AGN within this sample. Our modelling does not fit AGN components to the photometry, so if galaxies contain a significant AGN component it is likely we will overestimate the stellar mass. In particular, obscured AGN with red colors (e.g. Kocevski et al., 2023; Labbé et al., 2023) can mimic strong Balmer breaks, leading to inference of an aged stellar population and high stellar mass. Known as LRDs, a number of studies have found numerous candidates up to (see e.g. Greene et al., 2023; Furtak et al., 2023, 2023; Barro et al., 2023; Matthee et al., 2023; Labbe et al., 2023; Kokorev et al., 2024). Pérez-González et al. (2024) uses MIRI observations of LRDs to suggest that the characteristic photometry is driven mostly by stellar emission, and that LRDs are mostly compact and highly obscured starburst galaxies with young stellar ages. Other studies have struggled to distinguish between AGN and stellar models (Barro et al., 2023) or favour an AGN model (Noboriguchi et al., 2023). Further spectroscopic or MIRI observations can break the degeneracy between AGN and old stellar populations. We attempt to perform additional SED fitting using Prospector by including an AGN component, which models AGN emission lines for the most massive galaxies in our sample (M). In almost all cases the AGN fraction is less than 10%, and we see no reduction in stellar mass due to the inclusion of AGN component. The addition of an AGN model does not statistically improve the fit for these galaxies, although we can not strongly exclude the presence of AGN in some of our candidates as the AGN model within Prospector models only obscured AGN (Leja et al., 2018). Further fitting with a wider range of AGN models, which is outside the scope of this paper, is required to constrain the potential contribution of AGN emission to the observed photometry. Given the uncertainty around the true nature of these sources, we choose not to remove them from our sample. We discuss the impact of LRDs on the GSMF in Appendix A.

### v3

@@ 4.2 Prospector
In addition to our Bagpipes SED-fitting we also fit our sample using the Prospector package (Johnson et al., 2021) in order to compare the results of these two commonly used SED fitting tools. Prospector allows greater flexibility and control of model parameters than Bagpipes at the expense of computational time. Prospector uses Bayesian inference to determine galaxy stellar population properties and star formation histories. Prospector is built on Flexible Stellar Population Synthesis (FSPS, Conroy & Gunn, 2010)), using python-fsps (Johnson et al., 2023) and the Modules for Experiments in Stellar Astrophysics Isochrones and Stellar Tracks (MIST) stellar isochrones (Choi et al., 2016; Dotter, 2016). Prospector allows for very flexible star formation histories, including the use of non-parametric SFHs, which have become increasingly popular in the JWST era. We generally follow the prescription of Tacchella et al. (2022) for our Prospector model. We test both a traditional parametric star formation history as well as the “continuity bursty” non-parametric SFH used in Bagpipes. We allow the stellar mass to vary between 6 12 with a uniform logarithmic prior. For our parametric star formation history we use the ‘delayed-’ model, where .
@@ 4.3 Stellar Mass Comparisons
In this section we compare the consistency of the derived stellar masses between our fiducial and alternative Bagpipes models. Figure 1 shows a comparison of the derived galaxy properties of our fiducial Bagpipes run for all sources in our sample to our other Bagpipes results. The details of the models compared are shown in § 4.1. As discussed in § 4.1, we compare models with different priors, SPS models and SFH parametrizations. The top plot shows the systematic mass offset found when varying the dust law and priors, while the second row shows the same offset for parametric and non-parametric SFHs, as well as a BPASS SPS model (Stanway & Eldridge, 2018). For each Bagpipes model we show the Locally Weighted Scatterplot Smoothing fit (LOWESS, Cleveland, 1979). The LOWESS estimator is a non-parametric fitter for noisy data, and shows the overall trend between the mass discrepancy and fiducial stellar mass.
In Figure 2 and Figure 3 we show examples of the photometry, best-fitting Bagpipes & Prospector SEDs, and posterior redshift and stellar mass estimates for a selection of high-mass and or discrepant galaxies in our sample. We allow the redshift to vary within the EAZY-py posterior PDF for each Bagpipes fit, resulting in only small changes in redshift between individual Bagpipes results. 99.6% of photo- estimates fall within 15% of our fiducial Bagpipes redshift. The variation in stellar mass and other parameters between models will also depend on the available photometry, and in future surveys with more NIRCam medium bands or deep MIRI data different systematics may be observed. For this reason we place the detailed comparison of each model to the fiducial model, along with further discussion of other galaxy properties in Appendix A and in this section we summarise the impact each model has on the resultant stellar masses. A comparison between the stellar masses derived with Bagpipes and Prospector can also be found in the appendix.
Focusing first on the the upper panel of Figure 1, the dust extinction and metallicity priors appear to have little systematic effect on the stellar masses, although there is significant scatter seen in a few individual galaxies. Replacing the Calzetti et al. (2000) dust law with Charlot & Fall (2000) or Salim et al. (2018), we see a significant increase in stellar mass for the most massive galaxies, with the more flexible dust law typically favouring a steeper attenuation power-law than used in Calzetti et al. (2000). For these massive galaxies, when comparing to the default Calzetti et al. (2000) assumption, Salim et al. (2018) typically results in a better fitting SED with a lower , whereas the opposite is true for the Charlot & Fall (2000).
When we test alternative star formation history models, we see a larger stellar mass discrepancy when comparing to the non-parametric flexible SFH than when we compare to an alternative parametric SFH model (delayed exponential). In particular with the “continuity bursty” SFH we see offsets of dex with the same goodness of fit, as well as systematically larger stellar masses across the galaxy sample. A good example of this is seen in JADES-Deep-GS:9075 shown in Figure 3, where the observed flat photometry is consistent with a large range of possible star formation histories, the inferred masses of which can vary by dex with little variation in the . Exchanging the default BC03 SPS model for the BPASS model appears to have a complex effect on the derived stellar masses, with the magnitude and direction of the observed scatter dependent on the fiducial stellar mass.
@@ 4.4 Top-heavy IMF
In Figure 4 we show the results of modifying the IMF on the stellar mass estimates. We compare the masses derived with a standard Kroupa (2001) IMF to the two modified IMF models, finding significant decreases in stellar mass with very little change in the quality of the fits. Noting that due to the computational intensity of Prospector fitting we only fit a subset of our full sample, and calculate the median decrease in mass for both SFH models and HOT IMFs. For the sample where we use the HOT 60K IMF we include our full parent sample, however for the model we fit only galaxies with either or fiducial Bagpipes stellar mass . In terms of numbers, there are 221 galaxies in the HOT 45K IMF group, and 21 galaxies in the HOT IMF 60K group. Given that this IMF model is predicted to be applicable in the region of (Steinhardt et al., 2023), and the lack of mass-dependence in our results, we do not expect the excluded lower mass galaxies at to significantly impact our findings.
For the non-parametric “delayed” SFH model, we find a median decrease in stellar mass of 0.33 dex for the (HOT 45K IMF) galaxy sample. For the sample, we find a median decrease of 0.46 dex. For the “continuity bursty” SFH model, we find a median stellar mass decrease of 0.26 dex for the sample (HOT 45K IMF) and a decrease of 0.36 dex for the sample (HOT 60K IMF). As expected, we see a larger decrease in stellar mass for the HOT 60K model, which is more ‘top-heavy’ than the HOT 45K model.
@@ 5 Galaxy Stellar Mass Functions
We make a further cut to our galaxy sample here, removing the fields of El Gordo, SMACS-0723 and Clio. The depths and available filters of these datasets mean that they do not contribute significant volume to our GSMF estimates, but increase the redshift uncertainties (Adams et al., 2023). This reduces the number of galaxies used in the GSMF to 1092, and the total area used to 175 square arcminutes.
@@
Other distribution function estimators can been used that are more statistically robust that 1/, but 1/ is very commonly used in estimates of the UV LF and GSMF (e.g. Adams et al., 2023; Bhatawdekar et al., 2019; Navarro-Carrera et al., 2023; Stefanon et al., 2021; Weibel et al., 2024; Weaver et al., 2022; McLeod et al., 2023), so we choose to use it in order to compare our results directly with the literature. In Adams et al. (2023) we compare 1/, to the Lynden-Bell C-method, which is more statistically robust, when estimating the UV LF and find results consistent within 0.5 dex (Lynden-Bell, 1971; Woodroofe, 1985).
@@
where are the co-moving distances at the maximum/minimum detectable redshifts (capped at bin edges). are the survey areas, which are given in Table 1.
@@ 5.1 Detection Completeness
The magnitude range was swept in steps of 0.2, with 1000 sources inserted into the image at a time in order to prevent introducing artificial overdensities. We assume a size - luminosity relation of (Grazian et al., 2012), with reference size set to pc at , with MUV dependent intrinsic scatter modelled with a log-normal distribution calibrated from the results of Yang et al. (2022). The inclination angle of these sources was assumed to be uniformly random. We sweep a redshift range of 6.5 to 13.5 in steps of 0.5 to account for the changes in angular diameter distance and obtain a 2D dependence of completeness on redshift and absolute magnitude.
@@ 5.2 Selection Completeness and Contamination
We generate mock observations from these catalogues by estimating 1 errors in the measurements based on the average depths of our JWST reductions in each field (see Table 1). We then perturb the catalogue photometry in flux space within a Gaussian centered on the catalogue measurement and width equal to the 1 error in that filter. These perturbed fluxes are then run through our full selection procedure to measure how well our pipeline recovers the true redshifts of the sources. This allows us to compare the robustness of our selection criteria given the differing depths and filters in each set of observations.
@@ 5.2.3 Active Galactic Nuclei
Given the apparent prevalence of high- AGN detected by multi-wavelength and spectroscopic studies it is likely there is a population of galaxies containing AGN within this sample. Our modelling does not fit AGN components to the photometry, so if galaxies contain a significant AGN component it is likely we will overestimate the stellar mass. In particular, obscured AGN with red colors (e.g. Kocevski et al., 2023; Labbé et al., 2023) can mimic strong Balmer breaks, leading to inference of an aged stellar population and high stellar mass. Known as LRDs, a number of studies have found numerous candidates up to (see e.g. Greene et al., 2023; Furtak et al., 2023, 2023; Barro et al., 2023; Matthee et al., 2023; Labbe et al., 2023; Kokorev et al., 2024). Different selection techniques have been proposed for LRDs and it is not clear that all the objects grouped as LRDs with a given criteria indeed have the same origin. The term originated to categorise spectroscopically identified obscured broad-line AGN at high redshift, but has become more generally used to describe photometrically selected compact red galaxies, which may or may not be dominated by AGN emission. Photometric selections typically rely on a rest-optical color cut (F277W-F444W for ) as well as a compactness criteria, with spectroscopic analysis from Greene et al. (2023) suggesting that a F277W-F444W break exceeding 1.6 magnitudes identifies LRDs with an 80% AGN purity, without requiring an additional compactness criteria. The criteria of Labbé et al. (2023) and Kokorev et al. (2024) requires a weaker color selection, which may be selecting a mix of AGN and dusty star-forming systems. Williams et al. (2024) have used MIRI observations from 5 - 25m of a sample of likely AGN-dominated LRDs with strong breaks, finding a flattening of the red colors (in Fν) at longer wavelengths, which they suggest as evidence for obscured AGN which lack toroidal hot dust. In contrast Pérez-González et al. (2024) uses MIRI observations of a wider population of LRDs, which may not be a single class of object, to suggest that the characteristic photometry is driven mostly by stellar emission, and that LRDs are mostly compact and highly obscured starburst galaxies with young stellar ages. Other studies have struggled to distinguish between AGN and stellar models (Barro et al., 2023) or favour an AGN model (Noboriguchi et al., 2023). Stacking deep X-Ray observations of LRDs have also revealed a surprising lack of X-Ray emission Yue et al. (2024); Ananna et al. (2024); Maiolino et al. (2024) which is somewhat puzzling to explain via the AGN scenario, although high obscuration or intrinsic X-ray weakness in the accretion disc have been proposed (Maiolino et al., 2024). Further spectroscopic or MIRI observations can break the degeneracy between AGN and old stellar populations.
Whilst it is essentially impossible to distinguish between the contribution of AGN and stellar emission to an SED with only rest-optical photometry we attempt to perform additional SED fitting using Prospector by including an AGN component, which models AGN emission lines for the most massive galaxies in our sample (M). In almost all cases the AGN fraction is less than 10%, and we see no reduction in stellar mass due to the inclusion of AGN component. The addition of an AGN model does not statistically improve the fit for these galaxies, although we can not strongly exclude the presence of AGN in some of our candidates as the AGN model within Prospector models only obscured AGN (Leja et al., 2018). Further fitting with a wider range of AGN models, which is outside the scope of this paper, is required to constrain the potential contribution of AGN emission to the observed photometry. Given the uncertainty around the true nature of these sources, we choose not to remove them from our sample. We discuss the impact of LRDs on the GSMF in Appendix B.